OpenStudy (anonymous):

Which of the following sets could be the sides of a right triangle? A. {2,3, sqrt of 13} B. {2,2,4} C. (1,2, sqrt of 3 wouldn't it be B because it has all even numbers?

OpenStudy (chrisdbest):

No, C

OpenStudy (mathstudent55):

If a triangle is a right triangle, and the lengths of its sides are a, b, and c, as shown in the figure below, then this equation is true: \(a^2 + b^2 = c^2\) 

OpenStudy (chrisdbest):

When You do B: 2^2 + 2^2 = 8, and sqrt of 8 is not 4

OpenStudy (anonymous):

-sighs- im dumb...

OpenStudy (solomonzelman):

B is not a triangle, because 2+2\(\bcancel{\Large ~<~}4\)

OpenStudy (solomonzelman):

I meant the other way

OpenStudy (solomonzelman):

2+2 is not greater than 4

OpenStudy (mathstudent55):

In a right triangle, the sides labeled a and b are always the sides that form the right angle. They are called legs. The side labeled c is the hypotenuse. It is opposite the right angle, and it is always the longest side in a right triangle.

OpenStudy (mathstudent55):

For each choice do this: 1. Square each of the shorter sides and add the squares. 2. Then square the longest side. 3. If the sum in step 1. does not equal the number in step 2., then the triangle is not a right triangle.

OpenStudy (mathstudent55):

Here is choice B.: 1. \(2^2 + 2^2 = 4 + 4 = 8\) 2. \(4^2 = 16\) 3. Since 8 is not equal to 16, choice B is not a right triangle.

OpenStudy (mathstudent55):

Now do the same to choices A. and B. and see if they are right triangles.